A diffraction grating has 500 lines per mm. Its slit spacing or grating element will be equal to[A]. 500 mm [B]. 5 x 10-3 mm[C]. 2 x 10-5 mm [D]. 2 x 10-3 mm

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
Grating element = 1/N
 = 1/500
 = 0.2 * 10-2 mm
 = 2 * 10-3 mm
 
Hence option D is correct.
 
Regards
Arun (askIITians forum expert)